# API Pack: A Massive Multi-Programming Language Dataset for API Call Generation

### v3

@@ Abstract
We introduce API Pack, a massive multi-programming language dataset containing more than 1 million instruction-API call pairs to improve the API call generation capabilities of large language models. By fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack, we achieved around 10% and 5% higher accuracy compared to GPT-3.5 and GPT-4, respectively, in generating unseen API calls. Fine-tuning on API Pack enables cross-programming language generalization by leveraging a large amount of data in one language and small amounts of data from other languages. Scaling the training data to 1 million instances further improves the model’s generalization to new APIs not encountered during training. We open-source the API Pack dataset, trained models, and associated source code at https://github.com/zguo0525/API-Pack to facilitate further research.
@@ 1 Introduction
Large language models (LLMs) have shown promise in assisting with software engineering tasks [1, 2, 3, 4, 5, 6, 7, 8, 9], with a primary focus on code generation [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Our work builds on these advances but targets at the time-consuming task developers often face in finding application programming interface (API) call code examples. While finding an appropriate API is relatively easy with search, the challenge lies in generating the specific API call code. This involves understanding the API’s parameters, data types, and required operations, which can be complex and vary significantly across different APIs. Currently, developers typically seek out examples in documentations or API hubs [27], sifting through lengthy pages to locate relevant information [28]. This process is recognized as cumbersome and inefficient, as it requires developers to manually parse and adapt the example code to their specific use case. Our research aims to transform this workflow by using LLMs to identify appropriate API endpoints and generate corresponding API calls based on natural language instructions, thus automating the process of understanding and adapting API call code examples.
To achieve this research goal, we created API Pack, a dataset designed to improve the API call generation capabilities of LLMs. This massive multi-programming language dataset, containing over 1 million instances across 10 programming languages, is by far the largest open-source instruction dataset (see Table 1) for API call generation and API call intent detection. We evaluated the effectiveness of API Pack in improving API call generation through various experiments (see Section 4), and find that CodeLlama-13B, fine-tuned with only 20,000 Python instances from API Pack, surpasses the performance of GPT-3.5 and GPT-4 in generating API calls for entirely new APIs.
API Pack distinguishes itself from prior works [29, 30, 31] in two key aspects: multilinguality and scale. API Pack includes API calls across 10 different programming languages (see Table 1), allowing for a unique assessment of cross-lingual skill transfer - the extent to which improvements in one language apply to others. This is crucial for the practical application of LLMs in API call generation but has been underexplored in previous studies. Furthermore, with over a million instances spanning a wider range of real-world APIs and use cases compared to its predecessors, API Pack enables a more rigorous evaluation of generalization capabilities by controlling the volume of training data.
@@
Multi-lingual API calls?
@@
Data generation method: custom self-instruct self-instruct custom custom custom custom
@@
    Cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages.
@@
    Scaling the instruction data to 1 million instances improves generalization to new APIs, confirming the benefits of a larger dataset.
The paper is organized as follows: Section 2 reviews related works. Section 3 describes the construction of API Pack dataset. Section 4 outlines the experimental design for fine-tuning models on API Pack and evaluating their performance. The key findings are presented in Section 5. We conclude the paper in Section 6, and address the limitations in Section 7. API Pack dataset is licensed under the Creative Commons Attribution 4.0 International (CC BY 4.0) License. The associated source code is licensed under an MIT License.
@@ 2.1 Methods to Generate Instruction Data with LLMs
Manually creating instruction data is labor-intensive [32]. To automate this process, researchers have proposed using LLMs, despite the risk of generating low-quality instances. Two popular methods are Self-Instruct [33] and Evol-Instruct [32], and both methods involve instance generation and filtering. Self-Instruct mutates seed instances using in-context examples and filters them based on ROUGE-L similarity and other heuristics. Evol-Instruct generates instances via targeted prompting and filters them using predefined heuristics. Quality checks are crucial when fine-tuning models with synthetic data. Expert annotators can determine instance correctness [33], or strong LLMs like ChatGPT can score or classify instances [34] based on heuristics derived from a small sample. This approach has also been used to assess instruction complexity [35, 36].
@@ 2.2 LLMs for API Call Code Generation and Intent Detection
LLMs for code generation primarily focus on general coding tasks, evaluated by benchmarks like MBPP [37], HumanEval [38], and its variants [39, 40, 41]. Recent research explores LLMs and APIs in two areas: API call intent detection and API call code generation. API call intent detection identifies the appropriate API endpoint for a natural language task. LLMs for this purpose [31, 42, 43, 44, 45] work in hybrid architectures, where the LLM finds the API endpoint(s) and other components generate the code. These studies explore single and multi-API intent scenarios, with few achieving good results in the latter [31]. Our work focuses on creating LLMs that generate code to invoke API functionality, which is less explored than API call intent detection. Gorilla [30] generates API calls for loading pre-trained machine learning models from known model hubs. ToolBench [29] benchmarks tool manipulation capabilities of open-source LLMs. Another line of work explores function calling in language models [46, 47, 48], including ChatGPT. These models require pre-defined functions before execution, differing from our scenario where developers rely on the language model to guide them to the appropriate function or API based on natural language input.
@@ 3 API Pack
API Pack is an instruction dataset with more than one million instances. Each instance contains an input-output pair plus additional information about the API and respective endpoints. The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API to be used. The outputs are API call examples, specifically HTTP request code snippets curated from OpenAPI specification (OAS) files. The data in API Pack is sourced from four hubs that store OAS files: RapidAPI [^1] https://rapidapi.com/categories, APIGurus [^2] https://apis.guru/, Swaggerhub [^3] https://app.swaggerhub.com/search, and IBM’s public API Hub [^4] https://developer.ibm.com/apis/. Table 2 summarizes the total number of APIs, unique endpoints, and total instances (in different programming languages) contained within API Pack.
@@ 3.1 Data Pre-processing
The data pre-processing stage involves two main steps. First, we filter out OAS files with non-English data to maintain language consistency. We also remove OAS files with zero endpoints, as they lack the necessary information to generate meaningful API call examples or instructions. Next, we extract relevant information from the remaining OAS files. At the endpoint level, we collect the name, functionality, description, method, and path. At the API level, we gather the API name, description, and provider. To further ensure data quality, we apply a second filter to remove instances with missing crucial data for generating API calls (e.g., method, path, or endpoint name) and instructions (e.g., functionality, description, or endpoint name). This step guarantees that all instances in the dataset contain the necessary information for generating useful API call examples and instructions.
@@ 3.2 Create API Databases
We build an API database (DB) using the pre-processed data, with each instance containing endpoint information, API details, an API call example in a specific programming language, and a language identifier. We use OpenAPI Snippet [^5] https://www.npmjs.com/package/openapi-snippet to generate API calls in 10 programming languages (cURL, libcurl, Java, Node.js, Python, Go, Ruby, PHP, Swift, JavaScript) for endpoints from RapidAPI, APIGurus, and OpenAPI, while directly extracting API calls from OAS files for IBM API Hub. Appendix 9.5 provides details on the API DB instance structure and the programming language diversity for each source.
@@ 3.3 Instructions Generation
The instruction generation process consists of two main steps: creating high-quality instruction examples and generating instruction candidates. This two-step process is essential to ensure the generation of precise, clear, and useful instructions.
First, we create high-quality instruction examples for each API DB file. We randomly select three endpoints from each API and use their information (e.g., functionality, description, endpoint_name, path) along with the respective API name to fill in a list of predefined instruction templates. This process yields three instruction examples per API DB file. For the API Gurus and IBM API Hub sources, three authors manually review and refine the generated instruction examples by correcting grammatical errors, removing unnecessary information, and ensuring that the API name is present in all examples. During this manual review, we identify common error patterns in the generated instructions. To streamline the refinement process for the Swaggerhub and RapidAPI sources, we replace human labor with an LLM (Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2). We create a prompt (Prompt 1 in Appendix 9.8) based on the common error patterns identified during the manual review of the API Gurus and IBM API Hub instruction examples. This prompt is used to guide the LLM in refining the instruction examples for the Swaggerhub and RapidAPI sources.
Next, we use the high-quality instruction examples to generate instruction candidates for each instance in the API DBs. We prompt an LLM (Mistral-7B-Instruct-v0.2) with the endpoint’s information as input and provide the high-quality instruction examples as in-context examples (Prompt 2 in Appendix 9.8). The LLM generates five instruction candidates for each instance in the API DBs. Figure 2 shows an example of a generated instruction candidate, while Appendix 9.7 presents all five candidates generated for the same instance.
{minted}
[fontsize=, breaklines, breakanywhere] bash "instruction_candidates": [ "idx": 1, "candidate": "I’d like to confirm if 3D Secure is supported for a specific card brand or BIN using the Adyen BinLookup API. For 3D Secure 2, it will also provide device fingerprinting keys.", "input_tokens_mean": -0.5497539341557909, , … ], "api_call": "curl –request POST  –url https://paltest.adyen.com/pal/servlet/BinLookup/v40/get3dsAvailability  –header ’Authorization: Basic REPLACE_BASIC_AUTH’  –header ’content-type: application/json’  –data ’"additionalData":"property1":"string","property2":"string", "brands":["string"], "cardNumber":"string", "merchantAccount":"string", "recurringDetailReference":"string", "shopperReference":"string"’" …
Figure 2: Fragment of an API Pack instance in cURL.
@@ 3.4 Data Validation
To validate the generated instructions for each API call, we randomly select a sample of 121 instances, each containing one API call and five instructions. These instances have unique functionality, description, and endpoint_name. One of the authors manually labels the 605 instructions (121 instances x 5 instructions) as good or bad. The analysis of the instructions labeled as bad reveals three common characteristics: 1) containing multiple instructions instead of a single one, b) including unnecessary text before or after the main instruction, and c) failing to accurately use the correct API name. Based on these characteristics, we create three prompts with a fixed set of in-context examples to automatically label instructions. To select the best prompt, we use an LLM (Mistral-7B-Instruct-v0.2) to label the 605 manually classified instructions and compare the results with human-labeled results. We choose the prompt with the best performance (Prompt 4 in Appendix 9.8) and instruct Mistral-7B-Instruct-v0.2 to classify all instructions in our dataset as good or bad. Instances with fewer than two good instructions are removed from the dataset.
To select the best candidate from the subset of good instructions, we calculate the likelihood of an LLM re-generating the input text used to generate the instruction. We prompt Mistral-7B-Instruct-v0.2 with each instruction candidate (Prompt 3 in Appendix 9.8) and obtain the log probability of each token for the re-generated input text. We calculate the mean of these log probabilities (input_tokens_mean). Appendix 9.7 shows five instruction candidates and their respective input_tokens_mean. We select instructions labeled as good with the best input_token_mean for fine-tuning. Our final dataset contains 1,128,599 instances, each with a valid API call example and at least two high-quality instructions. Appendix 9.1 shows the data instances filtered out at each stage of the pipeline.
@@ 4 Experiments and Evaluation Framework
To optimize the instruction-following capabilities of the language models, we post-processed API Pack into two instruction-tuning templates: zero-shot and three-shot. The zero-shot template targets scenarios where the output is expected to be a straightforward inference from the given input. The three-shot template emphasizes the model’s ability to learn and generate output with in-context learning. The three-shot template is available in Appendix 5.
@@ 4.1 Experimental Settings
A. Selecting the baseline: Our first experimental setting serves the purpose of selecting a base model for the rest of our experiments. We fine-tune Mistral 7b [49], CodeLlama 7b and 13b, as well as Llama 2 13b [50] on a subset of API Pack (20,000 instances in Python programming language). We evaluate the performance of each resulting model fine-tuned on the 20k subset and use the base model of the best-performing fine-tuned model in the rest of our experiments.
@@
  • 3-shot retrieved & re-ranked: Five retrieved API examples, with three selected using a re-ranker model.
Note that these prompt settings are for testing/inference and differ from the instruction-tuning templates used in fine-tuning. We use bge-large-en-v1.5 [51] as the embedding model for retrieval and beg-reranker-large [52] for re-ranking. Appendix 9.4 illustrates the inference pipelines (0-shot, 3-shot) used to evaluate the models’ performance.
@@
D. Scaling experiment: We conduct a scaling experiment to investigate whether more API data improves a model’s generalization ability for unseen API. We fine-tune models on progressively larger API datasets with unique API calls: 10k, 20k, 40k, 80k, 100k, and 1 million instances. Our hypothesis is that exposure to a greater scale and diversity of APIs during fine-tuning will improve the model’s ability to generalize to new, unseen APIs.

### v4

@@ Abstract
We introduce API Pack, a massive multi-programming language dataset containing more than 1 million instruction-API call pairs to improve the API call generation capabilities of large language models. By fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack, we enable it to outperform GPT-3.5 and GPT-4 in generating unseen API calls. Fine-tuning on API Pack also facilitates cross-programming language generalization by leveraging a large amount of data in one language and small amounts of data from other languages. Scaling the training data to 1 million instances further improves the model’s ability to generalize to new APIs not used in training. To facilitate further research, we open-source the API Pack dataset, trained model, and associated source code at https://github.com/zguo0525/API-Pack.
@@ 1 Introduction
Large language models (LLMs) have shown promise in assisting with software engineering tasks [1, 2, 3, 4, 5, 6, 7, 8, 9], with a primary focus on code generation [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Our work builds on these advances but targets a time-consuming task developers often face in finding application programming interface (API) call code examples. While finding an appropriate API is relatively easy with search, the challenge lies in generating the specific API call code. This involves understanding the API’s parameters, data types, and required operations, which can be complex and vary significantly across different APIs. Currently, developers typically seek out examples in documentations or API hubs [27], sifting through lengthy pages to locate relevant information [28]. This process which requires developers to manually look for code examples for their specific use case is often cumbersome and tedious. Our work aims to transform this workflow by using LLMs to identify appropriate API endpoints and generate corresponding API calls based on natural language instructions, thus automating the process of searching the appropriate API call code examples.
To achieve this goal, we create API Pack, a dataset designed to improve the API call generation capabilities of LLMs. This multi-programming language dataset, containing over 1 million instances across 10 programming languages, is by far the largest open-source instruction dataset (see Table 1) for single API call generation and API call intent detection. API Pack distinguishes itself from prior works [29, 30, 31] in two key aspects: programming language diversity and dataset size. API Pack includes API calls across 10 different programming languages (see Table 1), allowing for a unique assessment of cross-lingual skill transfer - the extent to which improvements in one language apply to others. This important aspect has been under explored in previous studies even though it is crucial for the practical application of LLMs to API call generation use cases. Furthermore, with more than a million instances that span a wider range of real-world APIs and use cases compared to its predecessors (see Table 1), API Pack enables a more rigorous evaluation of generalization capabilities by controlling the volume of training data.
@@
Multi-programming language? ✓(10) ✗(Python) ✓(Curl, Python)
@@
Data generation method custom self-instruct self-instruct custom custom custom custom
@@
    Cross-lingual API call generation can be facilitated by fine-tuning models on a large amount of data in one language along with small amounts of data from other languages.
@@
    Scaling the instruction data from smaller instances to 1 million instances improves generalization to new APIs, validating the benefits of a larger dataset.
The paper is organized as follows: Section 2 reviews related works. Section 3 describes the construction of API Pack dataset. Section 4 outlines the experimental design for fine-tuning models on API Pack and evaluating their performance. The key findings are presented in Section 5. We conclude the paper in Section 6, and address limitations and future work plans in Section 7. We discuss ethical considerations related to API Pack’s use in Appendix 9.1. API Pack dataset is under the Creative Commons Attribution 4.0 International License. The associated source code is under an MIT License.
@@ 2.1 Methods to Generate Instruction Data with LLMs
Manually creating instruction data is labor intensive [32]. To automate this process, researchers have proposed the use of LLMs. Two popular methods to generate data via an LLM are Self-Instruct [33] and Evol-Instruct [32], both methods involve instance (instruction - response pairs) generation and filtering. Self-Instruct mutates seed instances using in-context examples and filters them based on ROUGE-L similarity and other heuristics. Evol-Instruct generates instances via targeted prompting and filters them using predefined heuristics. Custom LLM-based pipelines to generate synthetic data, such as the one we followed to create API Pack, also exist. In the case of code datasets, the pipelines typically involve mining high-quality code and using an LLM to generate instructions for each code data sample (e.g., [30], [29]). Post-quality filtering is crucial when using LLM-based synthetic data for fine-tuning. Although human annotators can determine the correctness of the instance [33], this kind of revision takes time and is thus not reasonable for large datasets. Therefore, recent work has proposed the use of strong LLMs (i.e., GPT-4 and Mistral AI models families) to score or classify instances [34] based on heuristics and examples derived from a small sample of data reviewed by human annotators. A similar approach has also been used to assess instruction complexity [35, 36].
@@ 2.2 LLMs for API Call Code Generation and Intent Detection
LLMs for code generation primarily focus on general coding tasks, evaluated by benchmarks like MBPP [37], HumanEval [38], and its variants [39, 40, 41]. Recent research explores LLMs and APIs in two areas: API call intent detection and API call code generation. API call intent detection identifies the appropriate API endpoint for a natural language task. LLMs for this purpose [31, 42, 43, 44, 45] work in hybrid architectures, where the LLM finds the API endpoint(s) and other components generate the code. These studies explore single and multi-API intent scenarios, with few achieving good results in the latter [31]. Our work focuses on training LLMs that generate code to invoke API functionality, which is less explored than API call intent detection. Gorilla [30] generates API calls for loading pre-trained machine learning models from known model hubs. ToolBench [29] benchmarks tool manipulation capabilities of open-source LLMs. Another line of work explores function calling in language models [46, 47, 48], including ChatGPT. These models require pre-defined functions before execution, differing from our scenario where developers rely on a language model to guide them to the appropriate function or API.
@@ 3 API Pack
API Pack is an instruction dataset with over one million instances. Each instance contains an input-output pair plus additional information about the API and respective endpoints. The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API. The outputs are API call examples, specifically HTTP request code snippets curated from OpenAPI specification (OAS) files. The data in API Pack is sourced from four hubs that store OAS files: RapidAPI [^1] https://rapidapi.com/categories, APIGurus [^2] https://apis.guru/, Swaggerhub [^3] https://app.swaggerhub.com/search, and IBM’s public API Hub [^4] https://developer.ibm.com/apis/. To the best of our knowledge, these API Hubs were the only sources with publicly available OAS files at a large scale at the time we collected data. From these sources, we kept all the APIs that passed our filtering criteria (see Section 3.1). Table 2 summarizes the total number of APIs, unique endpoints, and total instances in API Pack.
@@ 3.1 Data Pre-processing
This process involves two steps. First, as the presence of non-English metadata (e.g., API description, endpoint description) in the files was insignificant (less than 1%), we filter out OAS files with non-English data. We also remove OAS files with zero endpoints, as it is not possible to generate API calls for these APIs. Next, we extract relevant information from the remaining OAS files. We extracted information about the APIs, and about the endpoints that each API contains. At the API level, we gather the API name, description, and provider. At the endpoint level, we collect the name, functionality, description, method, and path. To further ensure data quality, we apply a second filter to remove instances with missing crucial data for generating API calls (e.g., method, path, or endpoint name) and instructions (e.g., functionality, description, or endpoint name). This step guarantees that all instances in the dataset contain the necessary information for generating useful API call examples and instructions.
@@ 3.2 Create API Databases
We build an API database (DB) using the pre-processed data, with each instance containing endpoint information, and API details. Then, we use OpenAPI Snippet [^5] https://www.npmjs.com/package/openapi-snippet library to generate API calls in 10 programming languages (cURL, libcurl, Java, Node.js, Python, Go, Ruby, PHP, Swift, JavaScript) for each API DB instance. We followed this process to create API calls with the data from RapidAPI, APIGurus, and OpenAPI. We directly extract API calls from OAS files for IBM API Hub as this source contained API calls in more than one programming language. Two factors influenced our selection of programming languages. First, the 10 programming languages included in API Pack appear in the StarCoder report of programming languages [23]. Second, given that we used an open-source library (OpenAPI Snippet) to generate API calls, language programming selection is inherently limited to the languages that this library supports. Appendix 9.6 provides details on the API DB instance structure and the programming language diversity for each source.
@@ 3.3 Instructions Generation
This process involves creating high-quality instruction examples, followed by generating instruction candidates based on these examples. First, we randomly select three endpoints from each API DB and use their information (e.g., functionality, description, endpoint_name, path) along with the respective API name to fill in a list of predefined instruction templates. This process yields three instruction examples per API DB file. For the API Gurus and IBM API Hub sources, three authors manually refine the generated instruction examples by correcting grammatical errors, removing unnecessary information, and ensuring that the API name is present in all examples. To streamline the refinement process of examples for the Swaggerhub and RapidAPI sources, we replace human labor with an LLM (Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2). We create a prompt to refine the instruction examples (Prompt 1 in Appendix 9.10) based on the heuristics identified during the manual review. We use this prompt to task the LLM with refining the instruction examples for the Swaggerhub and RapidAPI sources. Second, we use the high-quality instruction examples to generate five instruction candidates for each instance in the API DBs. We prompt an LLM (Mistral-7B-Instruct-v0.2) with the endpoint’s information as input and provide the high-quality instruction examples as in-context examples (see Prompt 2 in Appendix 9.10). Figure 7 in Appendix 9.8 shows an instance containing one instruction candidate along with its respective API call. Appendix 9.9 presents all five candidates generated for the same instance.
@@ 3.4 Data Validation
To validate the generated instructions for each API call, we randomly select a sample of 121 instances, each containing one API call and five instructions. These instances have unique functionality, description, and endpoint name. One of the authors manually annotates 605 instructions (121 instances x 5 instructions) as good or bad. The analysis of the instructions annotated as bad reveals three common characteristics: 1) containing multiple instructions instead of a single one, b) including unnecessary text before or after the main instruction, and c) failing to accurately use the correct API name. Based on these characteristics, we create three prompts with a fixed set of in-context examples to automatically annotate instructions via an LLM. To select the best prompt, we use an LLM (Mistral-7B-Instruct-v0.2) to annotate the 605 manually classified instructions and compare the results with human-annotated results. We choose the prompt with the best performance (Prompt 4 in Appendix 9.10) and instruct Mistral-7B-Instruct-v0.2 to classify all instructions in our dataset. Instances with fewer than two good instructions are removed from the dataset.
To select the best candidate from the subset of good instructions, we calculate the likelihood of an LLM re-generating the input text used to generate the instruction. We prompt Mistral-7B-Instruct-v0.2 with each instruction candidate (Prompt 3 in Appendix 9.10) and obtain the log probability of each token for the re-generated input text. We calculate the mean of these log probabilities (input_tokens_mean). Appendix 9.9 shows five instruction candidates and their respective input_tokens_mean. We select instructions annotated as good with the best input_token_mean for fine-tuning. Our final dataset contains 1,128,599 instances, each with a valid API call example and at least two high-quality instructions. Appendix 9.2 shows the data instances filtered out at each stage of the pipeline.
@@ 4 Experiments and Evaluation Framework
To optimize the instruction-following capabilities of the language models, we post-process API Pack into two instruction-tuning templates. The 0-shot-tuning template targets scenarios where the output is expected to be a straightforward inference from the given input. The 3-shot-tuning template, with random selections within the selected API, emphasizes the model’s ability to learn and generate output with in-context learning. The 3-shot-tuning template is available in Appendix 5.
@@ 4.1 Experimental Settings
A. Selecting a base model: Our first experimental study serves the purpose of selecting a base model for the rest of our experiments. We fine-tune Mistral 7b [49], CodeLlama 7b and 13b, as well as Llama 2 13b [50] on a subset of API Pack (20,000 instances in Python programming language). We evaluate the performance of resulting models and select the best-performing base model.
@@
  • 3-shot retrieved & re-ranked: Five retrieved API examples, with three selected using a re-ranker model in the final prompt.
The 3-shot random inference uses the same prompt template as 3-shot-tuning. We use bge-large-en-v1.5 [51] as the embedding model for retrieval and beg-reranker-large [52] for re-ranking. Appendix 9.5 illustrates the inference pipelines used to evaluate the models’ performance.
@@
D. Scaling experiment: We conduct a scaling experiment to investigate whether more API data improves a model’s generalization ability for unseen APIs. We fine-tune models on progressively larger API datasets with unique API calls on 10k, 20k, 40k, 80k, 100k, and 1 million instances. Our hypothesis is that exposure to a greater scale and diversity of APIs during fine-tuning will improve the model’s ability to generalize to new, unseen APIs.